OpenStudy (anonymous):

A hot bowl of soup cools according to Newton's law of cooling. Its temperature (in degrees Fahrenheit) at time t is given by T(t)=68+144e^(-.04t), where t is given in minutes. a. What was the initial temperature of the soup? b. What is the temperature of the soup after 15 minutes? c. How long after serving is the soup 125 degrees F?

OpenStudy (anonymous):

the initial temp is what you get when \(t=0\) namely \[68+144e^0=68+144\]

OpenStudy (anonymous):

that is 212 degrees, which is the boiling point of water

OpenStudy (anonymous):

to find the temp after 15 minutes, replace \(t\) by 15 and see what you get use a calculator

OpenStudy (anonymous):

is this correct? T(15)=68+144e^(-.04*15)=147.03 degrees

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=68%2B144e^%28-.04*15%29 that is what i get, yes

OpenStudy (anonymous):

for the last one , solve \[125=68+144e^{-.04t}\]for \(t\) do you know how to do that?

OpenStudy (anonymous):

good :)

OpenStudy (anonymous):

for c. is this correct? 125=68+144e^(-0.4t) (125-68)/144=0.395833333=e^(-.04t) -0.926762032=.04t t=23.169 After 23 minutes the soup is 125°F

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

let me check method looks right did you take the log at step 3?

OpenStudy (anonymous):

yes, it is right. looks like you got this under control right?

OpenStudy (anonymous):

thanks!! just double checking :)

OpenStudy (anonymous):

yw